OpenStudy (anonymous):

what are the x-intercepts for f(x)= x^2-6x+3 ? please help!!!

OpenStudy (anonymous):

when y:f(x)=0 u get the x intercepts so x^2 - 6x + 3 = 0 => x^2 - 6x + 9 = 6 => (x - 3)^2 = 6 => x - 3 = sqrt of 6 => x = \[3+\sqrt{6}\] and \[3-\sqrt{6}\]

OpenStudy (anonymous):

thanks so much! Thats what I got but i'm doing this online thing and it kept saying it was wrong, must be an error!